# Simultaneous All-Pay Auctions with Budget Constraints

### v1

@@ 4 Nash Equilibrium with Two Items
Given the profile $(B_{1},B_{2},v_{11},v_{12},v_{21},v_{22})$ with $B_{1}\neq B_{2}$, if $v_{s1}>L_{1}$ and $v_{s2}=L_{2}$, then the following strategies can form a Nash equilibrium:
$f_{s}(x_{s1},x_{s2})=\begin{cases}1-\frac{L_{2}}{v_{w2}},&x_{s1}=L_{1},\,x_{s2}=0;\\ \frac{1}{v_{w2}},&x_{s1}=L_{1},\,x_{s2}\in(0,T_{2}];\\ \frac{\frac{L_{1}}{v_{w1}}}{\sqrt{L_{1}^{2}+(L_{2}-T_{2})^{2}}},&\begin{array}[]{l}x_{s1}\in[0,L_{1}),\,x_{s2}\in(T_{2},L_{2}],\\ x_{s2}=-\frac{L_{2}-T_{2}}{L_{1}}x_{s1}+L_{2},\end{array}\end{cases}$
where $T_{2}=-\frac{L_{1}}{v_{w1}}v_{w2}+L_{2}$.
@@ B  Missing proofs in Section 4
Given the profile $(B_{1},B_{2},v_{11},v_{12},v_{21},v_{22})$ with $B_{1}\neq B_{2}$, if $v_{s1}=L_{1}$ and $v_{s2}>L_{2}$, then the following strategies can form a Nash equilibrium:
$f_{s}(x_{s1},x_{s2})=\begin{cases}\frac{1}{v_{w2}},&x_{s1}=0,\,x_{s2}\in[T_{2},L_{2}];\\ \frac{\frac{L_{1}}{v_{w1}}}{\sqrt{L_{1}^{2}+T_{2}^{2}}},&x_{s1}\in(0,L_{1}],\,x_{s2}\in[0,T_{2}),\,x_{s2}=-\frac{T_{2}}{T_{1}}x_{s1}+T_{2},\end{cases}$
where $T_{2}=\frac{L_{1}}{v_{w1}}v_{w2}$.
$f_{w}(x_{w1},x_{w2})=\begin{cases}\frac{1}{v_{s1}},&x_{w1}\in[T_{3},L_{1}],\,x_{w2}=0;\\ \frac{\frac{L_{2}}{v_{s2}}}{\sqrt{L_{2}^{2}+T_{3}^{2}}},&x_{w1}\in[0,T_{3}),\,x_{w2}\in(0,L_{2}],\end{cases}$

### v2

@@ 4 Nash Equilibrium with Two Items
Given the profile $(B_{1},B_{2},v_{11},v_{12},v_{21},v_{22})$ with $B_{1}\neq B_{2}$, if $v_{s1}>L_{1}$ and $v_{s2}=L_{2}$, then the following strategies form a Nash equilibrium:
$f_{s}(x_{s1},x_{s2})=\begin{cases}\frac{1}{v_{w2}},&x_{s1}=L_{1},\,x_{s2}\in(0,T_{2}];\\ \frac{\frac{L_{1}}{v_{w1}}}{\sqrt{L_{1}^{2}+(L_{2}-T_{2})^{2}}},&\begin{array}[]{l}x_{s1}\in[0,L_{1}),\,x_{s2}\in(T_{2},L_{2}],\\ x_{s2}=-\frac{L_{2}-T_{2}}{L_{1}}x_{s1}+L_{2},\end{array}\end{cases}$
where $T_{2}=-\frac{L_{1}}{v_{w1}}v_{w2}+L_{2}$. When $x_{s1}=L_{1},\,x_{s2}=0$, there is a mass point, the probability value is $1-\frac{L_{2}}{v_{w2}}$.
@@ B  Missing proofs in Section 4
Given the profile $(B_{1},B_{2},v_{11},v_{12},v_{21},v_{22})$ with $B_{1}\neq B_{2}$, if $v_{s1}=L_{1}$ and $v_{s2}>L_{2}$, then the following strategies form a Nash equilibrium:
$f_{s}(x_{s1},x_{s2})=\begin{cases}\frac{1}{v_{w1}},&x_{s1}=(0,T_{1}],\,x_{s2}=L_{2};\\ \frac{\frac{L_{2}}{v_{w2}}}{\sqrt{L_{2}^{2}+(L_{1}-T_{1})^{2}}},&\begin{array}[]{l}x_{s1}\in(T_{1},L_{1}],\,x_{s2}\in[0,L_{2}),\\ x_{s2}=\frac{L_{2}}{T_{1}-L_{1}}x_{s1}-\frac{L_{2}}{T_{1}-L_{1}}L_{1},\end{array}\end{cases}$
where $T_{1}=-\frac{L_{2}}{v_{w2}}v_{w1}+L_{1}$. When $x_{s1}=0,\,x_{s2}=L_{2}$, there is a mass point, the probability value is $1-\frac{L_{1}}{v_{w1}}$.
$f_{w}(x_{w1},x_{w2})=\begin{cases}\frac{1}{v_{s1}},&x_{w1}\in[T_{3},L_{1}],\,x_{w2}=0;\\ \frac{\frac{L_{2}}{v_{s2}}}{\sqrt{L_{2}^{2}+T_{3}^{2}}},&x_{w1}\in[0,T_{3}),\,x_{w2}\in(0,L_{2}],x_{w2}=-\frac{L_{2}}{T_{3}}x_{w1}+L_{2},\end{cases}$